# Leverage Cross-Attention for End-to-End Open-Vocabulary Panoptic Reconstruction

### v1

@@ Abstract
Index Terms:
Panoptic Reconstruction, Open-vocabulary segmentation, End-to-end, Cross-attention
@@ I Introduction
One intuitive but expensive method for open-vocabulary 3D panoptic reconstruction is to reconstruct and then segment 3D space by a 3D vision language model (VLM)[1, 2, 3, 4, 5], which requires a large amount of high-quality 3D annotated data. In contrast, 2D annotated data is far more accessible, enabling 2D open-vocabulary panoptic segmentation to reach near-application levels. Therefore, some feature-lifting methods  [6, 7, 8] propose to lift the capabilities of 2D VLMs to a 3D representation. 2D VLM features [9, 10, 11, 12] of images are distilled to a 3D visual language feature field by neural rendering [13] or back projection. While using the inner product of text prompts and distilled features enables flexible segmentation, its simple architecture results in imprecise masks.
@@
Our work is an extension version of a conference paper [19], where we introduce PanopticRecon, a two-stage zero-shot panoptic reconstruction method, that leverages the 2D VLMs and the geometric priors provided by graph inference and shows the competitive performance. In this paper, we have several extensions: 1) We propose a perspective of cross-attention to unify the previous methods and delve into new insight for architecture design; 2) We present PanopticRecon++, which injects instance prior and build panoptic supervision, and keeps the whole architecture end-to-end differentiable. 3) We conduct substantial experiments for 2D/3D segmentation, reconstruction, and rendering tasks on simulated and real-world indoor datasets.
@@ II-B Open-vocabulary Segmentation
2D Segmentation. Trained on massive image-text datasets, foundation models like CLIP [12] have achieved remarkable performance in aligning visual and textual representations. This lead to a surge of research exploring various zero-shot image tasks, such as object detection [35] and image captioning [36, 37], leveraging CLIP features. Building on this foundation, LSeg [10] and OpenSeg [9] extend the foundation models to semantic image segmentation. Then, Segment Anything Model (SAM) [11] further advances to segment arbitrary objects in any image, driven by user-provided or automatic generation prompts. This highlights the model’s ability to learn a generalizable notion of objects. More recently, Grounded SAM [38] innovatively combines object detection and segmentation foundation models to enable zero-shot 2D semantic and instance segmentation by text prompts.
@@ II-C Open-vocabulary Segmentation Reconstruction
Thanks to volume rendering-based reconstruction methods [13, 42, 43] supervised by images, it is possible to achieve open-vocabulary segmentation and reconstruction simultaneously leveraging 2D foundation models. LERF [7] and DFF [8] distill CLIP feature to a feature field to integrate language into 3D space, enabling flexible queries at any 3D point. However, these feature-lifting methods [7, 8, 18] still suffer from the inability to distinguish instance objects, rough segmentation boundaries, and high memory usage. When integrated with segmentation images from 2D foundation models, mask-lifting methods [19, 44, 45] can achieve more precise segmentation and reconstruction. GARField [44] lifts hierarchical 2D instance segmentation to 3D by contrastive learning. It optimizes the instance neural field such that pixels belonging to the same 2D mask are pulled closer and otherwise pushed apart, mitigating the problem of misaligned 2D instance mask IDs. Multi-scale segmentation reconstructions can be achieved, but this method requires post-processing to cluster features and may have difficulty in segmenting all instances well with a single parameter set. Gaussian Grouping [45] directly supervises 3D models with object-level 2D instance IDs from Grounded SAM [38] to ensure 3D object-level instance reconstruction. However, it faces the challenge of misaligned 2D instance IDs. Gaussian Grouping uses 2D tracking to align instance IDs, but tracking yields unstable performance. To align 2D instance IDs more precisely, we introduce geometric priors to understand 3D scenes comprehensively. In our previous work [19], we combined geometric priors with 3D graph inference to align 2D masks, which was achieved through a two-stage approach. In this work, we leverage the cross-attention between neural field and Gaussian-modulated tokens, combine the segmentation feature with the geometric prior, and propose an end-to-end framework for open-vocabulary panoptic reconstruction.
@@ IV-B Cross-Attention Design
where $P(\cdot|\cdot)$ represents the probability density of Gaussian distribution with specified mean and covariance. This ratio normalizes $\mathcal{G}$.
@@ IV-C Dynamic Token Adjustment
By calculating the IoM of 3D masks between each pair of tokens, we can identify duplicate tokens. We observe that duplicate tokens often exhibit an enclosing pattern, where a larger 3D instance mask encompasses multiple smaller 3D instance masks. So we keep only the token with the largest mask.
Split. Ideally, as the training of the scene model and instance tokens converges, a one-to-one correspondence between tokens and 3D instances should be established, indicating optimal assignments of each 2D image mask. However, if the initial token distribution is sparse, a single token may be assigned to multiple nearby objects, resulting in suboptimal assignments. Therefore, we monitor these assignments during training and dynamically introduce new tokens.
Specifically, given two 2D masks $M_{i}$ and $M_{j}$ from the same 2D instance segmentation image, if two masks both have the minimum cost with a token $Q_{i}$, but $Q_{i}$ is assigned to $M_{i}$, making $M_{j}$ assigned non-optimal token. Such result signals insufficiency of query tokens. If the signal keeps for iterations, we clone $Q_{i}^{{}^{\prime}}$ of $Q_{i}$ and append it to the instance tokens, which allows for assignment in further optimization e.g. the successful assignment between $Q_{i}^{{}^{\prime}}$ and $M_{j}$.
@@ V-B Panoptic Image Synthesis
Since the distribution is normalized to $1$, and all entries are larger than 00, we complete the validness of the label.
@@ V-D Implementation
Kimera - - - - - - - 69.13 85.12 - -
Panoptic NeRF 51.62 56.53 68.61 55.16 61.90· 72.66 81.13 55.42 66.43 71.75 90.50
Panoptic Lifting 66.59 67.19 86.36 63.80 72.52 63.73 70.01 72.44 84.40 81.50 71.42
Contrastive Lift 45.18 56.62 64.18 60.32 69.30 26.61 40.36 71.73 85.96 34.53 38.73
PVLFF 37.81 45.48 51.31 66.31 73.28 55.96 41.21 71.25 79.83 54.76 47.06
PanopticRecon 75.30 73.30 87.50 78.75 93.53 84.11 89.77 61.83 84.56 89.40 80.20
PanopticRecon++ 81.0 82.2 98.6 84.02 90.27 85.20 90.33 76.99 86.03 99.32 87.45
Kimera - - - - - - - 69.32 83.72 - -
Panoptic NeRF 52.12 58.68 75.81 51.99 62.04 49.29 61.97 52.27 68.53 70.42 83.09
Panoptic Lifting 60.24 65.80 91.11 73.10 84.39 43.00 61.65 56.25 77.47 53.80 58.10
Contrastive Lift 43.87 43.87 66.67 71.67 83.99 6.03 16.64 56.41 76.12 3.52 5.81
PVLFF 21.33 38.64 30.59 55.75 62.59 36.96 52.51 36.53 45.58 38.23 41.67
PanopticRecon 45.47 46.51 55.69 60.36 74.22 58.68 77.72 61.82 80.46 69.62 72.29
PanopticRecon++ 71.39 72.52 98.43 76.74 86.23 62.73 79.28 71.33 83.18 77.40 76.53
Kimera - - - - - - - 73.43 83.19 - -
Panoptic NeRF 61.11 64.34 84.66 56.39 70.93 62.47 68.59 53.15 69.54 73.80 89.13
Panoptic Lifting 58.06 63.44 91.85 66.28 75.65 41.11 47.54 72.62 86.36 38.21 40.20
Contrastive Lift 34.34 41.14 59.55 61.44 70.45 12.58 20.93 69.78 84.90 9.29 12.32
PVLFF 32.83 41.94 52.49 51.06 58.36 46.74 54.54 38.42 51.17 53.85 57.08
PanopticRecon 71.54 72.61 98.50 69.61 78.81 59.69 72.11 73.21 83.86 64.24 67.86
PanopticRecon++ 73.60 73.80 99.74 70.63 81.05 72.53 78.41 74.04 85.45 86.68 83.51
Kimera - - - - - - - 70.81 86.04 - -
Panoptic NeRF 67.89 78.19 86.00 72.95 88.58 69.43 76.38 61.40 79.00 73.46 89.85
Panoptic Lifting 46.65 51.41 71.90 68.45 81.79 35.68 60.51 67.51 83.59 38.33 51.17
Contrastive Lift 26.01 26.01 40.00 65.64 79.46 7.61 15.10 67.51 84.56 8.21 16.69
PVLFF 28.46 45.08 43.33 48.53 61.61 43.33 45.39 40.18 54.96 46.92 42.53
PanopticRecon 62.48 66.81 83.00 65.77 76.92 63.83 71.77 64.01 78.06 78.03 74.36
PanopticRecon++ 76.21 76.21 100 74.85 86.91 69.32 78.77 74.07 87.90 90.97 82.90
[…]
The table represents the quality on ”0087_02”, ”0088_00”, ”0420_01” and ”0628_02” four scenes from top to bottom.
@@
Kimera - - - - - - - 74.14 86.89 - -
Panoptic NeRF 51.97 62.18 84.69 52.04 68.96 56.68 64.48 52.76 74.91 72.85 88.47
Panoptic Lifting 63.55 69.25 83.03 80.04 88.63 53.69 68.91 76.04 87.59 53.17 59.05
Contrastive Lift 40.99 52.79 59.92 78.31 87.02 28.37 34.58 69.69 85.08 25.17 21.80
PVLFF 41.56 58.40 51.67 51.12 61.49 71.43 79.39 51.05 60.54 73.20 73.17
PanopticRecon 60.43 69.05 79.00 75.00 85.40 47.79 68.22 78.49 88.88 51.83 59.65
PanopticRecon++ 74.35 83.36 88.28 78.33 88.80 75.81 82.47 78.88 89.69 86.08 79.64
Kimera - - - - - - - 71.12 86.46 - -
Panoptic NeRF 30.24 36.32 42.36 41.78 70.98 23.08 56.87 39.73 69.65 68.17 83.65
Panoptic Lifting 68.87 78.56 86.96 77.69 86.40 44.46 57.84 72.92 86.56 56.49 64.22
Contrastive Lift 45.82 51.46 60.42 78.80 86.87 9.85 20.39 73.98 87.35 10.75 19.85
PVLFF 54.90 60.21 68.66 61.81 67.96 56.61 65.56 58.98 65.99 67.30 72.38
PanopticRecon 68.65 81.02 84.21 77.29 86.05 37.77 45.74 73.26 87.70 46.89 51.08
PanopticRecon++ 74.70 81.64 92.50 79.52 87.49 67.15 68.18 74.01 88.00 83.47 78.34
Kimera - - - - - - - 68.56 82.23 - -
Panoptic NeRF 56.19 63.99 88.03 56.56 73.47 48.64 48.44 48.62 67.13 67.59 82.04
Panoptic Lifting 81.02 84.64 94.44 86.29 93.97 70.35 78.78 70.06 83.12 85.94 75.87
Contrastive Lift 55.92 67.44 77.08 86.17 94.02 43.96 54.55 69.26 83.18 60.48 60.16
PVLFF 60.27 81.97 76.35 74.65 81.47 75.80 81.47 49.79 63.93 72.16 63.12
PanopticRecon 75.80 80.95 91.94 80.96 89.79 68.47 74.41 71.58 83.81 85.70 73.35
PanopticRecon++ 82.80 87.15 92.50 86.56 95.30 83.21 86.08 74.97 86.11 99.25 82.15
[…]
The table represents the quality on ”1ada7a0617”, ”5748ce6f01”, and ”f6659a3107” three scenes from top to bottom.
@@ VI-A Datasets
We evaluate our approach on three indoor datasets: one simulated dataset (Replica [52]) and two real-world datasets (Scannet-V2 [53] and Scannet++ [54]).
@@ VI-B Baseline
Then Panoptic Lifting (PL) [23], Contrastive Lift (CL) [21], PVLFF [18], and our previous work, PanopticRecon (PR) [19] are employed with the same problem setting to PanopticRecon++. They are representative works for panoptic reconstruction by lifting 2D observations, making them a fair comparison to PanopticRecon++. Both CL and PVLFF are two-stage methods based on feature clustering post-processing. PR, our previous work, also adopts a two-stage reconstruction pipeline. PL is a one-stage method but has no instance prior consideration.
@@ VI-C Evaluation Metrics
Rendering. We evaluate the color and depth of the rendered images separately. $PSNR$, $SSIM$, and $LPIPS$ are employed to comprehensively assess the quality of the rendered RGB images [57]. Absolute Relative Error ($AbsRel$ [58]) is utilized to reflect the overall discrepancy between the rendered depth map and the ground truth depth map.
@@ VI-D Experimental Setup
For a fair comparison, we consistently employ 2D semantic, instance, and panoptic segmentation images obtained through Grounded SAM as supervisory data for Kimera, PN, CL, PL, and PR. PN requires high-quality 3D object polygonal bounding box ground truth aligned with 2D instance labels. In our experiments datasets, such high-quality boxes are unavailable, so we provide them with object 3D rectangular bounding boxes obtained from instance mesh ground truth. For PVLFF, its default open-vocabulary segmentation methods are employed which is compatible with its hierarchical clustering.
We evaluate all frames of a scene but train on keyframes, ensuring each scene contains approximately 200-400 frames, to reduce computational costs and assess the results of novel views. The downsampling rate is adapted to the camera moving speed across scenes, but keep all comparative methods the same. All experiments were conducted on a single NVIDIA A100 GPU.
@@ VI-E Scene-level Instance Segmentation
As we can see, PanopticRecon++ demonstrates the best performance across all sequences and datasets, which owes to the cross-attention allowing for both spatial prior and direct back-propagation from instance loss. Compared with PN, which accesses the 3D ground truth, our methods also demonstrate better performance in most scenes, because PN is vulnerable to intersections between 3D rectangular bounding boxes in contrast to its original polygonal bounding boxes. When dealing with scenes where objects have less intersection, ”0628_02” scene, our method is slightly weaker. Since PanopticRecon++ has no access to the 3D ground truth, this comparison validates the superior performance.
@@
PR and PVLFF demonstrate a similar ranking on 3D instance segmentation as 2D instance segmentation. The only exception is ”5748ce6f01” scene, where the performance of PVLFF drops. As shown in Fig. 7, this drop is attributed to the presence of a large, ambiguous instance (a TV screen with its stand) in the ground truth, leading to ambiguity in segmentation. Since there are few observations of this ambiguous instance in the sequence, this issue is not reflected in the 2D instance segmentation metrics but clarified in 3D performance. For PL, the lack of instance prior explains the inferior performance in most sequences. These limitations are addressed by the 3D spatial prior, and again verify the performance of PanopticRecon++.

### v2

@@ I Introduction
One intuitive but expensive method for open-vocabulary 3D panoptic reconstruction is to reconstruct and then segment 3D space by a 3D vision language model (VLM)[1, 2, 3, 4, 5], which requires a large amount of high-quality 3D annotated data. In contrast, 2D annotated data is far more accessible, enabling 2D open-vocabulary panoptic segmentation to reach near-application levels. Therefore, some feature-lifting methods  [6, 7, 8, 9, 10] propose to lift the capabilities of 2D VLMs to a 3D representation. 2D VLM features [11, 12, 13, 14] of images are distilled to a 3D visual language feature field by neural rendering [15] or back projection. While using the inner product of text prompts and distilled features enables flexible segmentation, its simple architecture results in imprecise masks.
@@
Our work is an extension version of a conference paper [21], where we introduce PanopticRecon, a two-stage zero-shot panoptic reconstruction method. The most fundamental distinction between PanopticRecon++ and PanopticRecon lies in the shift to a fully end-to-end architecture. PanopticRecon leverages the 2D VLMs and the geometric priors provided by graph inference and shows competitive performance. In contrast, the key idea of PanopticRecon++ is to perform segmentation by introducing instance tokens modeled as 3D Gaussians. A cross-attention mechanism jointly leverages spatial priors and feature cues. Specifically, we have several extensions: 1) We propose a perspective of cross-attention to unify the previous methods and delve into new insight for architecture design; 2) We present PanopticRecon++, which injects instance prior and build panoptic supervision, and keeps the whole architecture end-to-end differentiable. 3) We conduct substantial experiments for 2D/3D segmentation, reconstruction, and rendering tasks on simulated and real-world indoor datasets.
@@ II-B Open-vocabulary Segmentation
2D Segmentation. Trained on massive image-text datasets, foundation models like CLIP [14] have achieved remarkable performance in aligning visual and textual representations. This lead to a surge of research exploring various zero-shot image tasks, such as object detection [37] and image captioning [38, 39], leveraging CLIP features. Building on this foundation, LSeg [12] and OpenSeg [11] extend the foundation models to semantic image segmentation. Then, Segment Anything Model (SAM) [13] further advances to segment arbitrary objects in any image, driven by user-provided or automatic generation prompts. This highlights the model’s ability to learn a generalizable notion of objects. More recently, Grounded SAM [40] innovatively combines object detection and segmentation foundation models to enable zero-shot 2D semantic and instance segmentation by text prompts. Additionally, several methods [41, 42, 43] perform tracking on 2D images to achieve consistent multi-frame 2D segmentation.
@@ II-C Open-vocabulary Segmentation Reconstruction
Thanks to volume rendering-based reconstruction methods [15, 47, 48] supervised by images, it is possible to achieve open-vocabulary segmentation and reconstruction simultaneously leveraging 2D foundation models. LERF [7] and DFF [8] distill CLIP feature to a feature field to integrate language into 3D space, enabling flexible queries at any 3D point. However, these feature-lifting methods [7, 8, 20] still suffer from the inability to distinguish instance objects, rough segmentation boundaries, and high memory usage. When integrated with segmentation images from 2D foundation models, mask-lifting methods [21, 49, 50] can achieve more precise segmentation and reconstruction. GARField [49] lifts hierarchical 2D instance segmentation to 3D by contrastive learning. It optimizes the instance neural field such that pixels belonging to the same 2D mask are pulled closer and otherwise pushed apart, mitigating the problem of misaligned 2D instance mask IDs. Multi-scale segmentation reconstructions can be achieved, but this method requires post-processing to cluster features and may have difficulty in segmenting all instances well with a single parameter set. Gaussian Grouping [50] directly supervises 3D models with object-level 2D instance IDs from Grounded SAM [40] to ensure 3D object-level instance reconstruction. However, it faces the challenge of misaligned 2D instance IDs. Gaussian Grouping uses a 2D tracker to align instance IDs, but tracking yields unstable performance. To align 2D instance IDs more precisely, we introduce geometric priors to understand 3D scenes comprehensively. In our previous work [21], we combined geometric priors with 3D graph inference to align 2D masks, which was achieved through a two-stage approach. In this work, we leverage the cross-attention between neural field and Gaussian-modulated tokens, combine the segmentation feature with the geometric prior, and propose an end-to-end framework for open-vocabulary panoptic reconstruction.
@@ IV-B Cross-Attention Design
where $P(\cdot|\cdot)$ denotes the Gaussian probability density with a specified mean and covariance, representing the spatial proximity to an instance. To normalize across different instance tokens, we divide $P(p_{k}|p_{q},\Sigma)$ by its peak $P(p_{q}|p_{q},\Sigma)$, yielding a normalized spatial prior $\mathcal{G}(p_{q},p_{k})\in[0,1]$, where higher values indicate stronger spatial affinity. This normalization ensures the prior reflects relative spatial compatibility rather than absolute density, making it suitable for modulating feature attention weights over the 3D scene.
@@ IV-C Dynamic Token Adjustment
which evaluates the 3D overlap between each pair of instance tokens. We observe that duplicate tokens often exhibit an enclosing pattern, where a larger 3D instance mask encompasses multiple smaller 3D instance masks. So we keep only the token with the largest mask.
Split. Ideally, as training converges, tokens form a one-to-one correspondence with 3D instances, yielding optimal assignments for 2D masks. When the initial token set is sparse, a single token may be the lowest-cost match for multiple nearby objects, causing persistent assignment conflicts. We address this by monitoring assignments and duplicating tokens on demand: Given masks $M_{i},M_{j}$ from the same image and a token $Q_{i}$, if the assignment cost function $\mathcal{C}(Q_{i},M_{i})=\min_{k}\mathcal{C}(Q_{k},M_{i})$ and $\mathcal{C}(Q_{i},M_{j})=\min_{k}\mathcal{C}(Q_{k},M_{j})$, yet the assignment assigns $Q_{i}$ to $M_{i}$ and forces $M_{j}$ to some $Q_{\ell}\neq Q_{i}$ with $\mathcal{C}(Q_{\ell},M_{j})>\mathcal{C}(Q_{i},M_{j})$ over several iterations, we clone $Q_{i}$ as $Q_{i}^{\prime}$ and append it to the token set, alleviating competition and enabling both $M_{i}$ and $M_{j}$ to obtain optimal matches.
@@ V-B Panoptic Image Synthesis
Since the distribution is normalized to $1$, and all entries are larger than 0, we complete the validness of the label.
@@ V-D Implementation
Kimera - - - - - - - 70.67 84.52 - -
Panoptic NeRF 58.19 64.44 78.77 59.12 70.86 63.46 72.02 55.56 70.88 72.36 88.14
Panoptic Lifting 57.89 61.96 85.31 67.91 78.59 45.88 59.93 67.21 82.96 52.96 55.22
Contrastive Lift 37.35 41.91 57.60 64.77 75.80 13.21 23.26 66.36 82.89 13.89 18.39
PVLFF 30.11 42.79 44.43 55.41 63.96 45.75 48.41 46.60 57.89 48.44 47.09
PanopticRecon 63.70 64.81 81.17 68.62 80.87 66.58 77.84 65.22 81.74 75.32 73.68
PanopticRecon++ 75.55 76.18 99.19 76.56 86.12 72.45 81.70 74.11 85.64 88.59 82.60
@@
Kimera - - - - - - - 71.27 85.19 - -
Panoptic NeRF 46.13 54.16 71.69 50.13 71.14 42.80 56.60 47.04 70.56 69.54 84.72
Panoptic Lifting 71.15 77.48 88.14 81.34 89.67 56.17 68.51 73.01 85.76 65.20 66.38
Contrastive Lift 47.58 57.23 65.81 81.09 89.30 27.39 36.51 70.98 85.20 32.13 33.94
PVLFF 52.24 66.86 65.56 62.53 70.31 67.95 75.47 53.27 63.49 70.89 69.56
PanopticRecon 68.29 77.01 85.00 77.75 87.08 51.34 62.79 74.44 86.80 61.47 61.36
PanopticRecon++ 77.28 84.05 91.09 81.47 90.53 75.39 78.91 75.95 87.93 89.60 80.04
@@ VI-A Datasets
We evaluate our approach on three indoor datasets and one outdoor dataset: one simulated dataset (Replica [57]) and three real-world datasets (Scannet-V2 [58] Scannet++ [59], and KITTI-360 [60]).
@@
KITTI-360. To evaluate the generalization capability of our method on larger-scale scenes, we further conduct experiments on the outdoor KITTI-360 dataset [60]. Specifically, we select a 200-meter segment from sequence 00 for evaluation, utilizing depth observations from a Velodyne HDL-64 LiDAR and a single perspective camera. Detailed experimental results and analysis are provided in the Appendix.
@@ VI-B Baseline
Then Panoptic Lifting (PL) [25], Contrastive Lift (CL) [23], PVLFF [20], and our previous work, PanopticRecon (PR) [21] are employed with the same problem setting to PanopticRecon++. They are representative works for panoptic reconstruction by lifting 2D observations, making them a fair comparison to PanopticRecon++ (PR++). Both CL and PVLFF are two-stage methods based on feature clustering post-processing. PR, our previous work, also adopts a two-stage reconstruction pipeline. PL is a one-stage method but has no instance prior consideration.
@@ VI-C Evaluation Metrics
Rendering. We evaluate the color and depth of the rendered images separately. $PSNR$, SSIMSSIM, and $LPIPS$ are employed to comprehensively assess the quality of the rendered RGB images [63]. Absolute Relative Error ($AbsRel$ [64]) is utilized to reflect the overall discrepancy between the rendered depth map and the ground truth depth map.
@@ VI-D Experimental Setup
In our experiments, we consistently adopt a fixed parameter setting for instance adjustment. In practice, we typically set the matching frequency threshold for redundancy pruning to 0 ane the $IoM=0.7$ for duplication pruning.
For the learning rate of instance tokens during training, we typically set the coordinate learning rate $l_{xyz}=2e{-3}$ and the scale learning rate $l_{s}=2e{-3}$. When the scene contains a large number of instances with significant variations in scale, we appropriately increase these two parameters to better adapt to the more complex instance distributions.
For a fair comparison, we consistently employ 2D semantic, instance, and panoptic segmentation images obtained through Grounded SAM as supervisory data for Kimera, PN, CL, PL, PR and PR++. Furthermore, to assess the impact of 2D VLM errors, we present an ablation study in the Appendix where we introduce different text prompts corresponding to objects that are unseen in the reconstructed scene.
PN requires high-quality 3D object polygonal bounding box ground truth aligned with 2D instance labels. On the three indoor datasets, such high-quality boxes are unavailable, so we provide them with object 3D rectangular bounding boxes obtained from instance mesh ground truth. For PVLFF, its default open-vocabulary segmentation methods are employed, which are compatible with its hierarchical clustering.
We evaluate all frames of a scene but train on keyframes, ensuring each scene contains approximately 200-400 frames, to reduce computational costs and assess the results of novel views. The downsampling rate is adapted to the camera moving speed across scenes, but keeps all comparative methods the same. All experiments were conducted on a single NVIDIA A6000 GPU. Concurrently, we also provide more detailed numerical results and analysis of the computational efficiency in the Appendix.
@@ VI-E Scene-level Instance Segmentation
As we can see, PanopticRecon++ demonstrates the best performance across all sequences and datasets, which owes to the cross-attention allowing for both spatial prior and direct back-propagation from instance loss. Compared with PN, which accesses the 3D ground truth, our methods also demonstrate better performance in most scenes, because PN is vulnerable to intersections between 3D rectangular bounding boxes in contrast to its original polygonal bounding boxes provided by KITTI-360. When dealing with scenes where objects have less intersection, ”0628_02” scene, our method is slightly weaker. Since PanopticRecon++ has no access to the 3D ground truth, this comparison validates the superior performance.
@@
PR and PVLFF demonstrate a similar ranking on 3D instance segmentation as 2D instance segmentation. The only exception is ”5748ce6f01” scene, where the performance of PVLFF drops. As shown in Fig. 7, this drop is attributed to the presence of a large, ambiguous instance (a TV screen with its stand) in the ground truth, leading to ambiguity in segmentation. Since there are few observations of this ambiguous instance in the sequence, this issue is not reflected in the 2D instance segmentation metrics but clarified in 3D performance. For PL, as shown in Tab. III and Tab. VI, the lack of spatial prior leads to multiple duplicate instance IDs assigned to objects that do not co-appear in the same frame, resulting in over-segmentation and inconsistent instance associations across views. This phenomenon directly contributes to its inferior instance segmentation performance.